# 00footnotetext: Shared first author; Shared last author

### v1

@@ Appendix I Model-based quality filters
In Appendix˜O we show how instruction tuning the above “model soup" for 80B additional tokens leads to strong performance on instruction tuning benchmarks and out-performs instruction tuned variants of similar 7B models such as Gemma-7B. In addition to the IT benchmarks covered in Appendix˜O, In Table˜29 we show that a small amount of instruction tuning provides large improvements in “Extended" evals at the cost of a small degradation in “Core" and “MMLU" evals. Notably we note that our GSM8k performance goes 2.5% to 52.5% which is comparable to other similar language models that mixed IT data into pretraining such as Gemma-7B.
@@
We show that the above strategy results in similar performance on regular evaluations as the starting checkpoint and significantly improves on the multi-document question-answering evaluation. We use the evaluation setup described in [93]: the context is filled with $k$ documents followed by a question. We ensure that one of the $k$ documents includes the answer to the question (a.k.a., golden document). We use $k=1,10,20,30$, and for each case, we run the evaluation multiple times by changing the position of the golden document in the context and report the average. Results are reported in table˜30. We demonstrate that long context adaptation results in a checkpoint (DCLM-8k) that matches the original model on regular evaluations and significantly improves multi-document QA showing its long-context capabilities.
@@
We note the compute cost of training runs for each competition scale in Table˜1. In total, we estimate that our runs for DCLM sum up to approximately 1M H100 hours. Our precise estimate from our experimental test bed is 772K H100 hours for training, but this is likely an underestimate due to additional compute that was not tracked, such as due to training failures.
@@
Appendix˜G discusses all downstream tasks we use for our evaluation. Below we mention them again, and specify their licenses.
@@
In addition to the above, as described in Sections˜5, 4.5 and 4.4 and Appendices˜L and O we make use of the following datasets:
@@
    StarCoder [88] and StarCoder2 [98] are used for some of our ablations (Section˜5. While constructed from permissive data by extracting datasets that mention permissive licenses (e.g. MIT, Apache 2.0), they involve various licenses, and as described in the Terms of Use [^8] https://huggingface.co/datasets/bigcode/the-stack#terms-of-use-for-the-stack, require the user to follow all terms-of-use and licenses of the different datasets it comprises of.
@@
    ProofPile2 [13] is used to scale up the dataset to the trillion tokens scale (Section˜5. The authors do not alter the licenses of underlying datasets and ask users to follow guidelines and licenses as described in these datasets.
@@
    GSM8k [39] was used in some of the ablations in Section˜5 and follow the MIT license.
@@
    RedPajama [156] is used for ablations in Section˜4.5 and Appendix˜L. Note that we do not release models or datasets that include this data. RedPajama filters the datasets it uses keeping only permissive licenses, and refers the user to adhere to underlying licenses where appropriate, as described in https://huggingface.co/datasets/togethercomputer/RedPajama-Data-1T.
@@
    transformers uses the Apache 2.0 License. [^9] https://github.com/huggingface/transformers/blob/main/LICENSE
@@
    • The purpose of DCLM and the associated DCLM-Pool and DCLM-baseline datasets is to enable the study of what makes a strong pretraining dataset for large language models. These models are transformative to society and act as the foundation of numerous applications, but they are often associated with steep costs. While prior work explores many curation techniques, it is often coupled with various architectural and training design choices and evaluated in different settings, making controlled comparison nearly impossible. This slows down progress and forces a lot of duplicate work between research teams. Prior work mainly focuses on data curation in the context of supervised datasets and smaller scales (see Section˜2 and Appendix˜B). In our initial release of DCLM, we focus on 53 downstream language understanding tasks that also include reasoning abilities, math, code, and more. For details see Section˜3.5 and Appendix˜G.
@@
    • Each instance represented a web-crawled page (document). It contains the URL and the corresponding HTML content. Each sample is also tagged with metadata about its crawl time and additional information such as the detected language, for processed instances such as those in DCLM-baseline. Additional information can be found in Appendix˜E.
@@
    • DCLM-Pool is an unfiltered web-text corpus comprised of all Common Crawl data prior to 2023. As such, it represent the full breadth of possible instances from this source. However, we note that Common Crawl does not cover the entire web data, due to reach and compute limitations for instance. For our DCLM-baseline, we use various filtering and deduplication strategies as described in Section˜4 in the explicit attempt to improve its quality for preatining, thus removing low-quality instances, and in doing so, becoming non-representative of the full set of instances. For a complete treatment and visualization of our data processing funnel, see Sections˜4, 4.3 and 4.2 and Appendix˜E.
@@
    • Each sample contains an a web-page url for and the extracted HTML content associated with. Additionally, each sample contains metadata fields shown in Table˜10 (e.g., WARC-Type, WARC-date, Content-Type etc.).
@@ 8. Q12
    • No. The evaluation procedure is made of tasks as described in Section˜3.5. We also attempt to prevent test set contamination in as described in Section˜4.6 and Appendix˜N.
@@ 10. Q14
    • Each sample is associated with a URL that links other external resources on the internet with no guarantee that the resources will exist in perpetuity or that that the resources will not change. However, the dataset itself contains already extracted HTML content and is thus self-contained for the purposes of this benchmark as described in Appendix˜C.
@@ 2. Q23
    • We begin by downloading the entire Common Crawl data prior to 2023. We ran Python-based processing scripts to parse these archives, filtering low-quality or irrelevant content, deduplicate samples and in some cases decontaminate against downstream tests sets, and compute various model-based features. We ran processes on hundreds of AWS CPU nodes for Common Crawl parsing and data downloading. Model-based features were run on GPU clusters. For software links see Q37 and Appendix˜R or refer to https://datacomp.ai/dclm.
@@ 1. Q39
    • The full dataset (and subsets) have been used to train hundreds of language models at various scales and compute budgets as presented in our main paper. We evaluate these models on our testbed of 53 zero- and few-shot downstream tasks. See Sections˜4 and 3.5.
@@ 4. Q48
    • We distribute our datasets in full, including extracted page content and associated metadata under a standard CC-BY-4.0 licence (see Appendix˜E). The code associated with DCLM is released under the MIT license. We also note that the use of this dataset is also subject to CommonCrawl’s Terms of Use as described in https://commoncrawl.org/terms-of-use.
@@ 4. Q55
    • Currently, there are no plans to update DCLM-Pool to maintain scientific integrity and comparability among participants in the DCLM competition. However, we will address user takedown requests (see Q56). DCLM-Pool is inherently noisy, and its release aims to encourage researchers to study dataset cleaning in the context of image-text samples.

### v2

@@ Appendix I Model-based quality filters
In Appendix O we show how instruction tuning the above “model soup" for 80B additional tokens leads to strong performance on instruction tuning benchmarks and out-performs instruction tuned variants of similar 7B models such as Gemma-7B. In addition to the IT benchmarks covered in Appendix O, In Table 29 we show that a small amount of instruction tuning provides large improvements in “Extended" evals at the cost of a small degradation in “Core" and “MMLU" evals. Notably we note that our GSM8k performance goes 2.5% to 52.5% which is comparable to other similar language models that mixed IT data into pretraining such as Gemma-7B.
@@
We show that the above strategy results in similar performance on regular evaluations as the starting checkpoint and significantly improves on the multi-document question-answering evaluation. We use the evaluation setup described in [95]: the context is filled with $k$ documents followed by a question. We ensure that one of the $k$ documents includes the answer to the question (a.k.a., golden document). We use $k=1,10,20,30$, and for each case, we run the evaluation multiple times by changing the position of the golden document in the context and report the average. Results are reported in table 30. We demonstrate that long context adaptation results in a checkpoint (DCLM-8k) that matches the original model on regular evaluations and significantly improves multi-document QA showing its long-context capabilities.
@@
We note the compute cost of training runs for each competition scale in Table 1. In total, we estimate that our runs for DCLM sum up to approximately 1M H100 hours. Our precise estimate from our experimental test bed is 772K H100 hours for training, but this is likely an underestimate due to additional compute that was not tracked, such as due to training failures.
@@
Appendix G discusses all downstream tasks we use for our evaluation. Below we mention them again, and specify their licenses.
@@
In addition to the above, as described in Sections 5, 4.5 and 4.4 and Appendices L and O we make use of the following datasets:
@@
    StarCoder [90] and StarCoder2 [101] are used for some of our ablations (Section 4). While constructed from permissive data by extracting datasets that mention permissive licenses (e.g. MIT, Apache 2.0), they involve various licenses, and as described in the Terms of Use [^8] https://huggingface.co/datasets/bigcode/the-stack#terms-of-use-for-the-stack, require the user to follow all terms-of-use and licenses of the different datasets it comprises of.
@@
    ProofPile2 [14] is used to scale up the dataset to the trillion tokens scale (Section 5). The authors do not alter the licenses of underlying datasets and ask users to follow guidelines and licenses as described in these datasets.
@@
    GSM8k [41] was used in some of the ablations in Section 4 and follows the MIT license.
@@
    RedPajama [160] is used for ablations in Section 4.5 and Appendix L. Note that we do not release models or datasets that include this data. RedPajama filters the datasets it uses keeping only permissive licenses, and refers the user to adhere to underlying licenses where appropriate, as described in https://huggingface.co/datasets/togethercomputer/RedPajama-Data-1T.
@@
    transformers uses the Apache 2.0 License.999 https://github.com/huggingface/transformers/blob/main/LICENSE
@@
    • The purpose of DCLM and the associated DCLM-Pool and DCLM-baseline datasets is to enable the study of what makes a strong pretraining dataset for large language models. These models are transformative to society and act as the foundation of numerous applications, but they are often associated with steep costs. While prior work explores many curation techniques, it is often coupled with various architectural and training design choices and evaluated in different settings, making controlled comparison nearly impossible. This slows down progress and forces a lot of duplicate work between research teams. Prior work mainly focuses on data curation in the context of supervised datasets and smaller scales (see Section 2 and Appendix B). In our initial release of DCLM, we focus on 53 downstream language understanding tasks that also include reasoning abilities, math, code, and more. For details see Section 3.5 and Appendix G.
@@
    • Each instance represented a web-crawled page (document). It contains the URL and the corresponding HTML content. Each sample is also tagged with metadata about its crawl time and additional information such as the detected language, for processed instances such as those in DCLM-baseline. Additional information can be found in Appendix E.
@@
    • DCLM-Pool is an unfiltered web-text corpus comprised of all Common Crawl data prior to 2023. As such, it represent the full breadth of possible instances from this source. However, we note that Common Crawl does not cover the entire web data, due to reach and compute limitations for instance. For our DCLM-baseline, we use various filtering and deduplication strategies as described in Section 4 in the explicit attempt to improve its quality for preatining, thus removing low-quality instances, and in doing so, becoming non-representative of the full set of instances. For a complete treatment and visualization of our data processing funnel, see Sections 4, 4.3 and 4.2 and Appendix E.
@@
    • Each sample contains a web-page url for and the extracted HTML content associated with. Additionally, each sample contains metadata fields shown in Table 10 (e.g., WARC-Type, WARC-date, Content-Type etc.).
@@ 8. Q12
    • No. The evaluation procedure is made of tasks as described in Section 3.5. We also attempt to prevent test set contamination in as described in Section 4.6 and Appendix N.
@@ 10. Q14
    • Each sample is associated with a URL that links other external resources on the internet with no guarantee that the resources will exist in perpetuity or that that the resources will not change. However, the dataset itself contains already extracted HTML content and is thus self-contained for the purposes of this benchmark as described in Appendix C.
@@ 2. Q23
    • We begin by downloading the entire Common Crawl data prior to 2023. We ran Python-based processing scripts to parse these archives, filtering low-quality or irrelevant content, deduplicate samples and in some cases decontaminate against downstream tests sets, and compute various model-based features. We ran processes on hundreds of AWS CPU nodes for Common Crawl parsing and data downloading. Model-based features were run on GPU clusters. For software links see Q37 and Appendix R or refer to https://datacomp.ai/dclm.
@@ 1. Q39
    • The full dataset (and subsets) have been used to train hundreds of language models at various scales and compute budgets as presented in our main paper. We evaluate these models on our testbed of 53 zero- and few-shot downstream tasks. See Sections 4 and 3.5.
@@ 4. Q48
    • We distribute our datasets in full, including extracted page content and associated metadata under a standard CC-BY-4.0 licence (see Appendix E). The code associated with DCLM is released under the MIT license. We also note that the use of this dataset is also subject to CommonCrawl’s Terms of Use as described in https://commoncrawl.org/terms-of-use.
@@ 4. Q55
    • Currently, there are no plans to update DCLM-Pool to maintain scientific integrity and comparability among participants in the DCLM competition. However, we will address user takedown requests (see Q56). DCLM-Pool is inherently noisy, and its release aims to encourage researchers to study dataset cleaning in the context of raw, web-crawled text samples.